# Steering MoE LLMs via Expert (De)Activation

### v1

@@ Abstract
Mixture-of-Experts (MoE) in Large Language Models (LLMs) routes each token through a subset of specialized Feed-Forward Networks (FFN), known as experts. We present SteerMoE, a framework for steering MoE models by detecting and controlling behavior-linked experts. Our detection method identifies experts with distinct activation patterns across paired inputs exhibiting contrasting behaviors. By selectively (de)activating such experts during inference, we control behaviors like faithfulness and safety without retraining or modifying weights. Across 11 benchmarks and 6 LLMs, our steering raises safety by up to +20% and faithfulness by +27%. In adversarial attack mode, it drops safety by -41% alone, and -100% when combined with existing jailbreak methods, bypassing all safety guardrails and exposing a new dimension of alignment faking hidden within experts.
Figure 1: Steering MoE models by routing through behavior-linked experts at inference enables lightweight, interpretable control without retraining or weight changes. Red and green FFNs are controlled by our method; others follow the router’s choice. Example outputs shown are generated by Qwen3-30B-A3B. (See more examples in Table 1)
@@ 1 Introduction
To this end, we introduce a general-purpose framework for steering MoE models by identifying behavior-linked experts through paired-example routing-difference detection. Our method compares expert activation patterns across prompt pairs exhibiting contrasting behaviors (e.g., safe vs. unsafe), and computes a simple risk difference score to quantify each expert’s behavioral association. At inference time, we then softly promote or suppress these experts by adjusting router logits, enabling lightweight behavioral steering without modifying model weights or requiring additional training. Our experiments span two critical dimensions of LLM behavior:
@@ 2.1 MoE Transformers Architectures
The layer then chooses the top-$k$ experts with the highest probabilities $\mathcal{T}=\operatorname{TopK}(\mathbf{p},k)$ and outputs the weighted mixture, so that only $k\ll E$ experts incur compute per token. [^1] The values $p_{i}$ will be renormalized to $\tilde{p}_{i}$ by dividing each by the total weight of the selected experts.
@@
  • GPT-OSS activates 4/32 and 4/128 in 20b and 120b models (OpenAI et al., 2025).
@@
  • DeepSeek-V2-Lite activates k=6k{=}6 of E=64E{=}64 experts per layer, which is 2B/16B parameters (DeepSeek-AI et al., 2024). (Omitted from experiments due to license restrictions.)
@@
  • Phi‑3.5‑MoE‑instruct activates 2/16 experts per layer and is 41.9B (Abdin et al., 2024).
@@ 2.2 Prior work on expert analysis
Beyond these in-paper snapshots, Lo et al. (2025) conducts a study over four public MoE LLMs. They observe that (i) neurons behave like finer-grained experts, (ii) routers favor experts with larger output norms, and (iii) expert diversity increases with depth, except for an outlier final layer.
@@ 3.1 Paired‑Example Routing‑Difference Detection
Let each example consist of a pair $(x^{(1)},x^{(2)})$, and let $\ell$ denote a specific layer in the MoE model. For every token $t$ in the input sequence and every expert $i\in{1,\dots,E}$, we track whether expert $i$ is among the top‑$k$ selected experts (i.e., routed to) at layer $\ell$ for token $t$. We define: $A^{(1)}_{i}$: the number of tokens in all $x^{(1)}$ for which expert $i$ is activated. $A^{(2)}_{i}$: the number of tokens in all $x^{(2)}$ for which expert $i$ is activated. $N^{(1)}$: the total number of tokens in all $x^{(1)}$ examples. $N^{(2)}$: the total number of tokens in all $x^{(2)}$ examples. From these, we compute the expert activation rates (i.e., empirical probabilities of activation):
$p^{(1)}_{i}=\frac{A^{(1)}_{i}}{N^{(1)}},\quad p^{(2)}_{i}=\frac{A^{(2)}_{i}}{N^{(2)}}.$ (3)
@@
$\Delta_{i}=p^{(1)}_{i}-p^{(2)}_{i}$, (4)
where $\Delta_{i}$ quantifies the difference in activation rate between the first and second prompt sets. A positive $\Delta_{i}$ indicates that expert $i$ is more frequently activated in the behavior shown in $x^{(1)}$, while a negative value suggests association with the behavior in $x^{(2)}$.
To steer the model we rank the experts by $|\Delta_{i}|$ (activation difference magnitude):
  • To promote the behavior associated with $x^{(1)}$, we activate experts with the most positive $\Delta_{i}$ and deactivate those with the most negative $\Delta_{i}$.
  • To promote the behavior associated with $x^{(2)}$, we apply the reverse: activate experts with the most negative $\Delta_{i}$ and deactivate those with the most positive $\Delta_{i}$.
@@ 3.2 Steering Setup
Let $\mathcal{A}^{+}\subseteq\{1,\dots,E\}$ and $\mathcal{A}^{-}\subseteq\{1,\dots,E\}$ denote the experts that must be activated or deactivated, respectively. Define $s_{\max}=\max_{j}s_{j}$ , $s_{\min}=\min_{j}s_{j}$ , and $\varepsilon>0\;(\text{e.g. }10^{-2})$
@@
For every $k\in\mathcal{A}^{+}$,
$s_{k}\leftarrow s_{\max}+\varepsilon$ (7)
@@
For every $k\in\mathcal{A}^{-}$,
$s_{k}\leftarrow s_{\min}-\varepsilon$ (8)
@@
where $\mathbf{h}$ is the input token representation to the MoE layer, and $\text{Expert}_{i}(\cdot)$ denotes the transformation performed by expert $i$. Note that, although the adjustments (Eq. 7, 8) set the target expert’s score to the maximum (or minimum) among all experts plus (or minus) a small constant $\varepsilon$, the modification remains minimal. This preserves the multi-expert structure of the weighted average in Eq.10, ensuring that multiple experts contribute meaningfully to the final output. In doing so, it avoids the extreme case where one expert is assigned probability 1, eliminating all other contributions, which would effectively collapse the mixture into a single-path model and undermine the intended specialization and redundancy benefits of MoE architectures. By contrast, our “soft” steering ensures that the selected expert is favored without fully suppressing others, enabling controlled behavior while maintaining overall model quality and stability.
@@ 4.1.2 Faithfulness Steering Results
Figure 2: Comparison of the off-the-shelf and steered models on faithfulness benchmarks. Accuracy is the proportion of examples in which the response remains faithful to the content of the provided document. The MC Test benchmark serves as a control dataset to ensure that the model’s general QA performance remains stable after steering. Modifying expert routing during inference improves performance on faithfulness benchmarks. (More models in Fig. A.2)
@@ 4.2.2 Safety Steering Results
Figure 3: Comparison of off-the-shelf and steered models on safety benchmarks shows that MoE steering via token routing can significantly adjust safety levels in both directions. Safety is measured by Llama-Guard-3-8B on harmful prompts, with ”Fluency” as a control to verify fluency. (More models in Fig. A.3)
@@
    Fluency: To evaluate fluency, we adopt an LLM-as-a-judge approach, leveraging the strong evaluative capabilities of recent language models (Liu et al., 2023; Wang et al., 2023). We use a 5-point Likert scale and report the normalized fluency score333(score1)/4(score-1)/4, as rated by Gemma-3-27B-it (Team et al., 2025) on the fluency of outputs in response to harmless prompts. (See the evaluation prompt in Table A.4.)
Figure 3 presents the results of the safety benchmarks. Across all models, steering expert routings toward unsafe behaviors degrades safety performance on the benchmarks. Notably, the “Fluency” benchmark indicates that this degradation occurs with minimal impact on the model’s overall fluency. Conversely, steering the model toward safe behaviors (safe experts) consistently improves safety, demonstrating the potential of expert-level interventions for alignment.
@@
Jailbreak Method
GPT-OSS
120b [Uncaptioned image]
Qwen3
Phi-3.5
OLMoE
@@
The experts most responsible for safety and RAG grounding cluster in the model’s middle layers (Fig. A.4 and A.5). This pattern echoes the findings of Muennighoff et al. (2025) and Jiang et al. (2024a), which attribute early and late layers mainly to vocabulary specialization. Together, these results suggest that high-level behavioral traits are shaped primarily in the model’s mid-depth.
@@ A.1.2 How Many Experts to (De)Activate?
Total Steer Faithful Steer Safe Steer Unsafe
@@
GPT-OSS-120B 4608 5 100 5 0 0 100
GPT-OSS-20B 768 10 50 5 0 0 20
Mixtral-8x7B-Instruct-v0.1 256 10 100 20 0 20 0
OLMoE-1B-7B-0125-Instruct 1024 0 50 5 0 10 125
Phi-3.5-MoE-instruct 512 10 75 5 0 5 50
Qwen3-30B-A3B 6144 0 500 15 0 5 480
@@
Table A.1 reports the number of manipulated experts for each model–task pair.
@@ A.2 Extra Figures and Tables
Expert
Activated
Expert
Deactivated
Total

### v2

@@ Abstract
Mixture-of-Experts (MoE) in Large Language Models (LLMs) routes each token through a subset of specialized Feed-Forward Networks (FFN), known as experts. We present SteerMoE, a framework to steer MoE models by detecting and controlling behavior-associated experts. We detect key experts by comparing how often they activate between paired inputs that demonstrate opposite behaviors (e.g., safe vs. unsafe). By selectively activating or deactivating such experts during inference, we control behaviors like faithfulness and safety without fine-tuning. Across 11 benchmarks and 6 LLMs, our steering raises safety by up to +20% and faithfulness by +27%. Alternatively, unsafe steering drops safety by -41% alone, and -100% when combined with existing jailbreak methods, bypassing all safety guardrails. Overall, SteerMoE offers a lightweight, effective, and widely applicable test-time control, while revealing unique vulnerabilities in MoE LLMs.
Figure 1: Steering MoE models by routing through behavior-linked experts at inference enables lightweight, interpretable control. Red and green FFNs are controlled by our method; others follow the router’s choice. Generations are from Qwen3-30B-A3B. (See more examples in Table 1)
@@ 1 Introduction
To this end, we introduce a general-purpose framework for steering MoE models by identifying behavior-linked experts. Our method compares expert activation rates between prompt pairs exhibiting contrasting behaviors (e.g., safe vs. unsafe), and computes a simple risk difference score to quantify each expert’s behavioral association. At inference time, we then promote or suppress these experts by adjusting router logits, enabling lightweight behavioral steering without modifying model weights or additional training. Our experiments span two critical dimensions of LLM behavior:
@@ 2.1 MoE Transformers Architectures
The layer then chooses the top-$k$ experts with the highest probabilities $\mathcal{T}=\operatorname{TopK}(\mathbf{p},k)$ and outputs the weighted mixture, so that only $k\ll E$ experts incur compute per token in each layer. [^1] The values $p_{i}$ will be renormalized to $\tilde{p}_{i}$ by dividing each by the total weight of the selected experts.
@@
  • GPT-OSS activates 4/32 and 4/128 experts in 20b and 120b models [^2] All $k/E$ counts are per layer; totals: $(k\cdot layer)/(E\cdot layers)$. (e.g., GPT-OSS-120B: 144/4608; Tab. A.2). (OpenAI et al., 2025).
@@
  • DeepSeek-V2-Lite activates k=6k{=}6 of E=64E{=}64 experts, which is 2B/16B parameters (DeepSeek-AI et al., 2024). (Omitted from experiments due to license restrictions.)
@@
  • Phi‑3.5‑MoE‑instruct activates 2/16 experts and is 41.9B (Abdin et al., 2024).
@@ 2.2 Prior work on expert analysis
Beyond these in-paper snapshots, Lo et al. (2025) conducts a study over four public MoE LLMs. They observe that (i) neurons behave like finer-grained experts, (ii) routers favor experts with larger output norms, and (iii) expert diversity increases with depth, except for an outlier final layer. Concurrent work on multilingual routing finds language-specific expert selection in early and late layers, with strong cross-lingual alignment in middle layers (Bandarkar et al., 2025).
@@ 3.1 Paired‑Example Routing‑Difference Detection
Let each example consist of a pair $(x^{(1)},x^{(2)})$, and let $\ell$ denote a specific layer in the MoE model. For every token $t$ in the input sequence and every expert $i\in{1,\dots,E}$, we track whether expert $i$ is among the top‑$k$ selected experts (i.e., routed to) at layer $\ell$ for token $t$. We define: $A^{(1)}_{\ell,i}$: the number of tokens in all $x^{(1)}$ for which expert $i$ in layer $\ell$ is activated. $A^{(2)}_{\ell,i}$: the number of tokens in all $x^{(2)}$ for which expert $i$ in layer $\ell$ is activated. $N^{(1)}$: the total number of tokens in all $x^{(1)}$ examples. $N^{(2)}$: the total number of tokens in all $x^{(2)}$ examples. From these, we compute the expert activation rates (i.e., empirical probabilities of activation):
$p^{(1)}_{\ell,i}=\frac{A^{(1)}_{\ell,i}}{N^{(1)}},\quad p^{(2)}_{\ell,i}=\frac{A^{(2)}_{\ell,i}}{N^{(2)}}.$ (3)
@@
$\Delta_{\ell,i}=p^{(1)}_{\ell,i}-p^{(2)}_{\ell,i}$, (4)
where $\Delta_{\ell,i}$ quantifies the difference in activation rate between the first and second prompt sets. A positive $\Delta_{\ell,i}$ indicates that expert $i$ in layer $\ell$ is more frequently activated in the behavior shown in $x^{(1)}$, while a negative value suggests association with the behavior in $x^{(2)}$.
To steer the model we rank the experts by $|\Delta_{\ell,i}|$ (activation difference magnitude):
  • To promote the behavior associated with $x^{(1)}$, we activate experts with the most positive $\Delta_{\ell,i}$ and deactivate those with the most negative $\Delta_{\ell,i}$.
  • To promote the behavior associated with $x^{(2)}$, we apply the reverse: activate experts with the most negative $\Delta_{\ell,i}$ and deactivate those with the most positive $\Delta_{\ell,i}$.
@@ 3.2 Steering Setup
Let $\mathcal{A}^{+}\subseteq\{1,\dots,E\}$ and $\mathcal{A}^{-}\subseteq\{1,\dots,E\}$ denote the experts that must be activated or deactivated in this layer. Define $s_{\max}=\max_{j}s_{j}$ , $s_{\min}=\min_{j}s_{j}$ , and $\varepsilon>0\;(\text{e.g. }10^{-2})$
@@
For every $e\in\mathcal{A}^{+}$,
$s_{e}\leftarrow s_{\max}+\varepsilon$ (7)
@@
For every $e\in\mathcal{A}^{-}$,
$s_{e}\leftarrow s_{\min}-\varepsilon$ (8)
@@
where $\mathbf{h}$ is the input token representation to the MoE layer, and $\text{Expert}_{i}(\cdot)$ denotes the transformation performed by expert $i$. Note that, although the adjustments (Eq. 7, 8) set the target expert’s score to the maximum (or minimum) among all experts plus (or minus) a small constant $\varepsilon$, the modification remains minimal. This preserves the multi-expert structure of the weighted average in Eq.10, ensuring that all top-$k$ experts contribute meaningfully to the final output. In particular, it avoids the extreme case where the target expert (or experts) in $\mathcal{A}^{+}$ receive a total probability of 1, distributed only among themselves, while the remaining experts in the top-$k$ set $\mathcal{T}$ receive zero probability. Such a collapse would effectively reduce the mixture to only a few active paths, undermining the benefits of MoE architectures and deviating from the behavior of the trained model. By contrast, our “soft” steering ensures that the selected experts are favored without fully suppressing others, enabling controlled behavior while maintaining overall model quality and stability.
@@ 4.1.2 Faithfulness Steering Results
Figure 2: Comparison of steered versus non-steered model performance on faithfulness benchmarks. Accuracy is the proportion of examples in which the response remains faithful to the content of the provided document. The MC Test benchmark serves as a control dataset to ensure that the model’s general QA performance remains stable after steering. Modifying expert routing during inference improves performance on faithfulness benchmarks. (More models in Fig. A.3)
@@ 4.2.2 Safety Steering Results
Figure 3: Comparison of steered versus non-steered model performance on safety benchmarks shows that MoE steering via token routing can significantly adjust safety levels in both directions. Safety is measured by Llama-Guard-3-8B on harmful prompts, with ”Fluency” as a control to verify fluency. (More models in Fig. A.4)
@@
    Fluency: To evaluate fluency, we adopt an LLM-as-a-judge approach, leveraging the strong evaluative capabilities of recent language models (Liu et al., 2023; Wang et al., 2023). We use a 5-point Likert scale and report the normalized fluency score444(score1)/4(score-1)/4, as rated by Gemma-3-27B-it (Team et al., 2025) on the fluency of outputs in response to harmless prompts. (See the evaluation prompt in Table A.7.)
Figure 3 presents the results of the safety benchmarks. Across all models, steering expert routings toward unsafe behaviors degrades safety performance on the benchmarks. Notably, the “Fluency” benchmark indicates that this degradation occurs with minimal impact on the model’s overall fluency. Conversely, steering the model toward safe behaviors (safe experts) generally improves safety, demonstrating the potential of expert-level interventions for alignment.
@@
Jailbreak Method GPT-OSS 120b [Uncaptioned image] Qwen3 [Uncaptioned image] Phi-3.5 [Uncaptioned image] OLMoE [Uncaptioned image]
@@
The experts most responsible for safety and RAG grounding cluster in the model’s middle layers (Fig. A.5 and A.6). This pattern echoes the findings of Muennighoff et al. (2025) and Jiang et al. (2024a), which attribute early and late layers mainly to vocabulary specialization. Together, these results suggest that high-level behavioral traits are shaped primarily in the model’s mid-depth.
@@ 5 Conclusions
Limitations
Our method relies on several assumptions. First, the approach requires access to models with a Mixture-of-Experts (MoE) architecture. While MoEs are increasingly common in large-scale systems, the technique cannot be directly applied to dense models without an analogous notion of per-token expert routing. Second, our method relies on paired inputs that exhibit clear contrasts in the targeted behavior. While such pairs are often easy to gather or synthesize for well-defined tasks, they may require additional curation or domain knowledge for more subtle or emergent behaviors. Third, determining how many experts to adjust depends on model-specific factors, including the number of experts, routing sparsity, and whether the model was trained with sufficiently strong incentives for sparse and stable expert utilization. The optimal configuration may therefore vary across architectures and tasks. We provide additional discussion and practical guidance on these considerations in the Appendix.
@@ A.1.1 Why Risk Difference?
Figure A.1: Pairplot of different scoring methods (Risk Difference, Log-Odds Ratio, and Paired t-test) for the detection of faithfulness related experts.
Score Method Before Steering 100 200 500 1000
Risk Difference 81% 86% 93% 97% 97%
Log-Odds Ratio 81% 84% 84% 82% 78%
Table A.1: Faithfulness on MQauke dataset using Qwen3 under deactivating top k experts detected by Risk Difference and Log-Odds Ratio.
Our preliminary analysis showed that RD is the best for steering. Figure A.1 empirically illustrates that the log-odds ratio exhibits high variance around zero, discussed before. Table A.1 compares downstream performance using each scoring method and shows that RD yields the strongest empirical results, aligning with the intuition we provide in the main text.
@@ A.1.2 How Many Experts to (De)Activate?
Active / Total Steer Faithful Steer Safe Steer Unsafe
@@
GPT-OSS-120B 144 / 4608 5 100 5 0 0 100
GPT-OSS-20B 96 /   768 10 50 5 0 0 20
Mixtral-8x7B-Instruct-v0.1 64 /   256 10 100 20 0 20 0
OLMoE-1B-7B-0125-Instruct 128 / 1024 0 50 5 0 10 125
Phi-3.5-MoE-instruct 64 /   512 10 75 5 0 5 50
Qwen3-30B-A3B 384 / 6144 0 500 15 0 5 480
@@
Table A.2 reports the number of manipulated experts for each model–task pair. Hyperparameter selection is an important part of our method. Different MoE models vary widely in the number of experts, the number of active experts per layer, and overall parameter counts. Models also differ in how sparsely behaviors are distributed across experts due to differences in pre-training paradigms (Muennighoff et al., 2025). As a result, it is natural and expected to observe variation in the number of experts identified across models and tasks. Crucially, once a model and task are fixed, the selected experts generalize consistently across all benchmarks for that task, as demonstrated in our results. In practice, we recommend a simple grid search over the number of activated/deactivated experts, jointly considering task performance and generation fluency (illustrated in Figure A.2).
@@ A.1.3 Why Deactivation Is Preferable to Activation
A.1.4 Prompted vs. Free Generation in Detection phase
In safety detection analyses of section 4.2 tokens after ”Assistant:” in already prompted examples are used for detection, but actual unsafe generation patterns of a model may differ during free generation. There are a few reasons why our approach of teacher-forcing remains appropriate compared with free-form generation: i) Feasibility: For certain behaviors, especially unsafe generations, free-form generation is often infeasible. Most models are explicitly trained to avoid generating harmful outputs, making it difficult to elicit such behavior naturally, even with extensive prompting. In these cases, teacher forcing becomes necessary. ii) Effectiveness: Despite using this method, our attack setup already achieves an Attack Success Rate gain from 0% to 100%, showing that it is effective in practice. iii) Generalizability of the Method: The use of prompted completions is a design choice, not a limitation of our method. If one has access to a paired dataset of free generations (e.g., unsafe vs. safe outputs) and the ability to annotate them, our method can be applied just as effectively to those. In this sense, SteerMoE is flexible and can operate over any paired dataset with behavior labels, making this adaptability a strength rather than a constraint.
A.1.5 Transfer Results on MultiJail
Jailbreak Method English Italian Thai
Direct Instruction 99.7% 100% 99.4%
AIM 100% 100% 100%
SteerMoE 94.3% 90.2% 87.9%
SteerMoE + AIM 9.5% 11.4% 7.3%
Table A.3: Safe response rates of GPT-OSS-120b on 315 MultiJail examples (lower = stronger attack). SteerMoE is competitive alone and yields the best results combined with AIM even on different languages.
We evaluate cross-lingual transfer of SteerMoE using the MultiJail dataset (Deng et al., 2023), which consists of 315 harmful prompts translated into multiple languages. Table A.3 reports the safe-response rates across languages. Notably, the results indicate that SteerMoE’s steering effects generalize cross-lingually: In this case, the same safety-related experts appear to be used to refuse harmful prompts in multiple languages, and deactivating those experts can jailbreak the model in other languages as well, leading to less than 12% safety, even though detection was run only on an English paired dataset.
A.1.6 Expert Selection Baselines
Method (Modifying 100 Experts) Safe Response Rate
Before Steering 100%
Random selection of experts (Seed 0) 100%
Random selection of experts (Seed 1) 100%
Random selection of experts (Seed 2) 100%
Bottom k Experts 100%
Steering with Our Selection 0%
Table A.4: Safe response rates of GPT-OSS-120b on 50 AdvBench examples. Naive selection baselines fail to steer the model in a meaningful way.
Figure A.4 reports an expanded set of baseline results for expert selection to activate or deactivate. All baselines show no effect on safety, confirming that only our method identifies behavior-relevant experts. As expected, these experts are sparse and task-specific, making naive selection ineffective for meaningful steering.
@@ A.2 Extra Figures and Tables
Expert Activated Expert Deactivated Total
@@
Algorithm 1 Computing Expert Risk Differences from Two Routing Distributions
1:Inputs:
2:   Two routing–probability tensors $\mathcal{P}^{(A)},\mathcal{P}^{(B)}\in\mathbb{R}^{T\times L\times E}$
3:   Number of routed experts $k$
4:Initialize:
5: Counters $\mathsf{Hit}^{(A)},\mathsf{Miss}^{(A)},\mathsf{Hit}^{(B)},\mathsf{Miss}^{(B)}\in\mathbb{N}^{L\times E}$ to zero.
6:(A) Count expert activations and non-activations
7:for each token $t=1\dots T$ do
8:  for each layer $\ell=1\dots L$ do
9:   Determine top-$k$ experts for $\mathcal{P}^{(A)}_{t,\ell,:}$.
10:   Increment $\mathsf{Hit}^{(A)}_{\ell,e}$ for each activated expert $e$.
11:   Increment $\mathsf{Miss}^{(A)}_{\ell,e}$ for all remaining experts.
12:   Repeat the same procedure for $\mathcal{P}^{(B)}$ using $\mathsf{Hit}^{(B)}$ and $\mathsf{Miss}^{(B)}$.
13:  end for
14:end for
15:(B) Compute activation frequencies and risk differences
16:for each layer $\ell$ and expert $e$ do
17:  Activation rates:
$r^{(A)}_{\ell,e}=\frac{\mathsf{Hit}^{(A)}_{\ell,e}}{\mathsf{Hit}^{(A)}_{\ell,e}+\mathsf{Miss}^{(A)}_{\ell,e}},\qquad r^{(B)}_{\ell,e}=\frac{\mathsf{Hit}^{(B)}_{\ell,e}}{\mathsf{Hit}^{(B)}_{\ell,e}+\mathsf{Miss}^{(B)}_{\ell,e}}$.
18:  Risk difference:
$\Delta_{\ell,e}=r^{(A)}_{\ell,e}-r^{(B)}_{\ell,e}$.
19:end for
20:Output: Risk differences $\Delta_{\ell,e}$ for all layers and experts.
Algorithm 2 Inference-Time Expert Steering in an MoE Layer
1:Inputs:
2:   Token representations $H\in\mathbb{R}^{N\times d}$
3:   Gating network $\mathcal{G}$ producing router logits over $E$ experts
4:   Steering vector $w\in\mathbb{R}^{E}$ for this layer
5:   Margin parameter $\varepsilon>0$ (e.g. $\varepsilon=0.01$)
6:Step 1: Compute base router scores
[…]
9:Step 2: Identify positively and negatively steered experts
[…]
12:Step 3: Clamp log-scores by per-token max/min
13:for each token index $i=1\dots N$ do
14:  $m^{\max}_{i}\leftarrow\max_{j}S_{i,j}$ {max log-score for token ii}
15:  $m^{\min}_{i}\leftarrow\min_{j}S_{i,j}$ {min log-score for token ii}
16:  for each expert $e\in P$ do
17:   $S_{i,e}\leftarrow m^{\max}_{i}+\varepsilon$
18:  end for
19:  for each expert $e\in N$ do
20:   $S_{i,e}\leftarrow m^{\min}_{i}-\varepsilon$
21:  end for
22:end for
23:Step 4: Route through experts with steered scores
24: Use $S$ as the routing scores to compute the MoE layer output.
25:Output: Steered mixture-of-experts output for all tokens.